solve the system. Assume that a and b represent nonzero constants.
ax+by=6
ax-3by=4
The x value of the solution is:

Answer by stanbon(75887)

solve the system. Assume that a and b represent nonzero constants.
ax+by=6
ax-3by=4
The x value of the solution is:
-------------------------
Subtract the 2nd equation from the 1std to get:
4by= 2
y = 1/2b
--------------
Substitute to solve for "x":
ax+b(1/2b) = 6
ax 1/2 + 6
ax = 11/2
x = 11/2a
================
Cheers,
Stan H.